OpenStudy (anonymous):

what is the rule with fractions that have exponents? For example, I am trying to evaluate 36 = (1/6)^x

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

(1/6)^x=(1^x)/(6^x)=1/(6^x)

OpenStudy (anonymous):

6^x=1/36 x=-2

OpenStudy (anonymous):

36=(1/6)^(x)=6^(-x) (because 1/(anything) is the same as (anything) to the power -1 (anything^-1) ) then you should know 36=6^2 so 2 = -x Leading to x=-2

OpenStudy (anonymous):

I understand now - thank you! I forgot that property!